# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

### v1

@@ Appendix E Evaluation results
anglic
english
eng_Latn eng_Latn 81.90 83.35 81.95 83.43
arabic
standard arabic
arb_Arab eng_Latn 65.57 73.83 65.57 73.85
najdi arabic
ars_Arab eng_Latn 59.42 69.02 59.14 68.94
ta’izzi-adeni arabic
acq_Arab eng_Latn 58.84 68.72 58.64 68.62
moroccan arabic
ary_Arab eng_Latn 54.65 58.30 54.61 58.14
egyptian arabic
arz_Arab eng_Latn 54.53 65.87 53.86 65.70
south levantine arabic
ajp_Arab eng_Latn 54.09 64.13 53.42 63.81
north mesopotamian arabic
acm_Arab eng_Latn 52.95 58.94 52.84 58.75
levantine arabic (a:north)
apc_Arab eng_Latn 52.14 61.74 51.40 61.31
tunisian arabic
aeb_Arab eng_Latn 47.54 50.18 47.42 50.20
common turkic
north azerbaijani
azj_Latn eng_Latn 59.76 73.17 59.20 73.17
central oghuz (m:spoken)
tur_Latn eng_Latn 59.14 74.47 58.37 74.52
south azerbaijani
azb_Arab eng_Latn 46.05 41.82 44.58 39.24
gallo-italian
venetian
vec_Latn eng_Latn 65.15 68.52 64.99 68.55
lombard
lmo_Latn eng_Latn 59.38 56.39 59.34 56.16
ligurian
lij_Latn eng_Latn 57.82 57.70 56.70 57.16
high german
luxemburgish
ltz_Latn eng_Latn 60.34 47.33 60.01 46.21
limburgan
lim_Latn eng_Latn 50.80 59.88 50.31 59.75
italian romance
italian
ita_Latn eng_Latn 73.71 78.08 73.71 78.19
sicilian
scn_Latn eng_Latn 62.69 56.17 62.66 55.82
kurdish
central kurdish
ckb_Arab eng_Latn 40.98 44.93 37.40 39.59
northern kurdish
kmr_Latn eng_Latn 39.10 63.45 33.93 63.26
latvian
latvian
lvs_Latn eng_Latn 60.14 73.57 59.95 73.63
east latvian
ltg_Latn eng_Latn 48.62 54.19 47.02 53.54
norwegian
norwegian bokmål (m:written)
nob_Latn eng_Latn 72.44 79.44 72.45 79.51
norwegian nynorsk (m:written)
nno_Latn eng_Latn 68.08 71.16 68.10 71.06
sinitic
classical-middle-modern sinitic (o:simplified)
zho_Hans eng_Latn 68.52 72.61 68.54 72.57
classical-middle-modern sinitic (o:traditional)
zho_Hant eng_Latn 61.72 64.89 61.48 64.49
cantonese
yue_Hant eng_Latn 60.44 68.02 60.27 67.41
sotho-tswana (s.30)
northern sotho
nso_Latn eng_Latn 39.38 40.18 35.06 35.98
southern sotho
sot_Latn eng_Latn 39.36 39.30 34.62 34.16
southwestern shifted romance
spanish
spa_Latn eng_Latn 75.13 79.00 75.15 79.09
portuguese (a:european)
por_Latn eng_Latn 73.73 79.18 73.73 79.22
galician
glg_Latn eng_Latn 73.39 78.48 73.39 78.55
occitan
oci_Latn eng_Latn 68.48 63.09 68.47 62.96
@@
cluster variety target-code count
Finetune (mBERT)
Finetune (XLM-R)
Zeroshot (mBERT)
Zeroshot (XLM-R)
ICL Mistral
anglic
english (a:scotland)
english–gbr 440
76.38
70.34
71.82
63.15
70.18
southern african english
english–zaf 440
76.66
71.18
71.49
63.87
71.14
new zealand english
english–nzl 440
76.71
71.39
71.22
63.69
70.95
australian english
english–aus 440
75.66
70.89
71.20
62.28
69.23
southeast american english
english–usa 440
77.26
71.50
71.17
63.71
71.76
irish english
english–irl 440
75.52
70.73
70.92
62.15
70.64
philippine english
english–phl 440
76.37
70.64
70.47
62.22
70.11
nigerian english
english–nga 440
73.61
68.33
69.10
61.27
68.10
indian english (a:north)
english–ind_n 440
74.62
68.03
68.84
61.25
68.99
kenyan english
english–kenya 440
72.59
66.68
68.72
58.64
64.91
indian english (a:south)
english–ind_s 440
71.93
66.88
66.49
60.36
65.13
arabic
arabic (a:bahrain)
arabic–bhr 921
77.52
72.11
53.25
53.28
55.33
arabic (a:jordan)
arabic–jor 921
77.35
71.29
52.72
53.72
54.93
arabic (a:saudi-arabia)
arabic–sau 921
77.88
72.11
52.72
53.24
55.66
algerian arabic
arabic–dza 921
77.85
72.34
52.56
53.52
55.18
tunisian arabic
arabic–tun 921
76.72
71.64
52.28
52.94
55.03
moroccan arabic
arabic–mar 921
76.73
71.57
51.86
52.17
53.92
egyptian arabic
arabic–egy 921
76.53
70.75
51.80
51.99
54.66
bengali
bengali (a:west bengal)
bengali–ind 113
68.62
73.27
32.30
36.39
50.06
bengali (a:dhaka)
bengali–dhaka 113
67.37
74.24
31.79
35.52
51.73
korean
seoul (m:spoken)
korean–korn 276
10.15
31.91
7.26
19.62
65.79
korean (a:south-eastern, m:spoken)
korean–kors 276
9.92
31.01
7.22
20.08
64.63
swahili
swahili (a:tanzania)
swahili–tanzania 472
63.54
62.30
38.24
39.38
55.70
swahili (a:kenya)
swahili–kenya 472
72.25
70.53
37.97
41.59
58.71
@@
cluster variety
dialect-code
support
precision (mBERT)
precision (XLM-R)
recall (mBERT)
recall (XLM-R)
F1 (mBERT)
F1 (XLM-R)
anglic
english (a:uk)
english:en-gb
249
98.10
89.57
83.13
71.59
90.00
79.58
north american english
english:en-us
349
93.27
88.14
83.38
82.09
88.05
85.01
arabic
aleppo
arabic:ale
200
59.50
58.50
59.50
48.35
59.50
52.94
algerian arabic
arabic:alg
272
79.00
65.50
58.09
62.68
66.95
64.06
arabian peninsula arabic (a:yemen)
arabic:san
177
60.50
55.50
68.36
56.63
64.19
56.06
egyptian arabic (a:alx)
arabic:alx
192
70.50
74.50
73.44
66.82
71.94
70.45
egyptian arabic (a:asw)
arabic:asw
221
56.00
52.00
50.68
45.02
53.21
48.26
egyptian arabic (a:cai)
arabic:cai
130
35.50
36.50
54.62
72.28
43.03
48.50
egyptian arabic (a:kha)
arabic:kha
244
63.50
55.50
52.05
44.05
57.21
49.12
fez. meknes
arabic:fes
196
60.00
59.50
61.22
56.40
60.61
57.91
gilit mesopotamian arabic
arabic:bag
203
57.50
47.50
56.65
49.48
57.07
48.47
gulf arabic (a:doh)
arabic:doh
205
50.00
43.50
48.78
45.55
49.38
44.50
gulf arabic (a:jed)
arabic:jed
196
58.00
46.00
59.18
40.89
58.59
43.29
gulf arabic (a:mus)
arabic:mus
178
38.50
49.50
43.26
42.67
40.74
45.83
gulf arabic (a:riy)
arabic:riy
311
62.00
56.50
39.87
37.92
48.53
45.38
levantine arabic (a:north-dam)
arabic:dam
148
37.50
24.50
50.68
42.98
43.10
31.21
libyan arabic (a:ben)
arabic:ben
238
56.50
47.50
47.48
52.78
51.60
50.00
north mesopotamian arabic (a:bas)
arabic:bas
186
49.50
39.00
53.23
49.68
51.30
43.70
north mesopotamian arabic (a:mos)
arabic:mos
188
71.50
70.00
76.06
69.31
73.71
69.65
rabat-casablanca arabic
arabic:rab
153
50.00
40.00
65.36
60.61
56.66
48.19
sfax
arabic:sfx
215
62.50
64.50
58.14
48.13
60.24
55.13
south levantine arabic (a:south-amm)
arabic:amm
177
40.50
35.00
45.76
35.53
42.97
35.26
south levantine arabic (a:south-jer)
arabic:jer
202
48.50
47.00
48.02
40.34
48.26
43.42
south levantine arabic (a:south-sal)
arabic:sal
167
46.00
66.50
55.09
59.11
50.14
62.59
standard arabic
arabic:msa
244
75.00
98.00
61.48
95.61
67.57
96.79
sunni beiruti arabic
arabic:bei
192
58.00
52.50
60.42
68.18
59.18
59.32
tripolitanian arabic
arabic:tri
201
66.00
60.00
65.67
60.30
65.84
60.15
tunisian arabic (a:tun)
arabic:tun
164
52.50
37.00
64.02
56.49
57.69
44.71
greek
cypriot greek (r:casual, m:written, i:other)
greek:cg_other
81
74.14
84.48
53.09
56.32
61.87
67.59
cypriot greek (r:casual, m:written, i:twitter)
greek:cg_twitter
36
51.11
44.44
63.89
68.97
56.79
54.05
modern greek (r:casual, m:written, i:twitter)
greek:smg_twitter
94
89.83
100.00
56.38
53.15
69.28
69.41
high german
central alemannic (a:be)
swiss-dialects:be
389
72.89
51.58
71.21
62.42
72.04
56.48
central alemannic (a:bs)
swiss-dialects:bs
340
73.50
57.83
75.88
61.14
74.67
59.44
central alemannic (a:lu)
swiss-dialects:lu
335
72.91
65.13
75.52
59.47
74.19
62.17
central alemannic (a:zh)
swiss-dialects:zh
359
78.84
73.33
75.77
63.73
77.27
68.19
sinitic
mandarin chinese (a:mainland, o:simplified)
mandarin_simplified:m
986
97.90
96.10
99.29
90.66
98.59
93.30
mandarin chinese (a:mainland, o:traditional, i:synthetic)
mandarin_traditional:m
977
96.80
92.10
99.08
95.74
97.93
93.88
mandarin chinese (a:taiwan, o:simplified)
mandarin_simplified:t
1014
99.30
90.10
97.93
95.85
98.61
92.89
mandarin chinese (a:taiwan, o:traditional, i:synthetic)
mandarin_traditional:t
1023
99.10
95.90
96.87
92.39
97.97
94.11
southwestern shifted romance
brazilian portuguese
portuguese:pt-br
627
96.94
92.35
90.91
84.98
93.83
88.51
latin american spanish
spanish:es-ar
207
81.06
9.25
88.89
91.30
84.79
16.80
portuguese (a:european)
portuguese:pt-pt
349
91.45
83.64
70.49
63.92
79.61
72.46
portuguese (m:written)
portuguese:pt
15
9.70
0.00
86.67
0.00
17.45
0.00
spanish
spanish:es
290
74.21
74.53
81.38
47.69
77.63
58.16
spanish (a:europe)
spanish:es-es
492
90.99
83.33
82.11
78.89
86.32
81.05
@@
arabic
standard arabic
arb_Arab arb_Arab 85.25 83.96 86.71 82.27
ta’izzi-adeni arabic
acq_Arab arb_Arab 84.96 82.05 86.44 81.98
najdi arabic
ars_Arab arb_Arab 84.80 84.39 87.41 83.33
north mesopotamian arabic
acm_Arab arb_Arab 82.97 80.95 84.77 80.36
south levantine arabic
ajp_Arab arb_Arab 81.82 80.16 84.16 79.05
levantine arabic (a:north)
apc_Arab arb_Arab 81.59 80.15 83.76 79.88
egyptian arabic
arz_Arab arb_Arab 81.02 76.38 84.43 81.03
tunisian arabic
aeb_Arab arb_Arab 79.45 72.88 83.97 77.33
moroccan arabic
ary_Arab arb_Arab 73.87 79.14 78.76 78.55
common turkic
north azerbaijani
azj_Latn azj_Latn 80.46 79.87 82.00 79.55
central oghuz (m:spoken)
tur_Latn azj_Latn 79.10 84.41 80.61 79.51
south azerbaijani
azb_Arab azj_Latn 65.90 67.08 69.71 68.37
gallo-italian
venetian
vec_Latn ita_Latn 76.72 70.68 75.07 74.28
lombard
lmo_Latn ita_Latn 69.92 59.90 70.65 64.56
ligurian
lij_Latn lij_Latn 66.81 63.42 74.03 57.78
high german
luxemburgish
ltz_Latn nld_Latn 74.74 58.50 77.86 64.83
limburgan
lim_Latn nld_Latn 71.09 65.83 71.12 65.73
italian romance
italian
ita_Latn ita_Latn 87.67 84.92 86.68 85.83
sicilian
scn_Latn ita_Latn 75.22 59.71 72.70 59.47
kurdish
northern kurdish
kmr_Latn ckb_Arab 33.23 68.21 10.45 5.71
central kurdish
ckb_Arab ckb_Arab 13.10 19.37 16.86 12.38
latvian
latvian
lvs_Latn lvs_Latn 76.35 83.75 80.63 82.80
east latvian
ltg_Latn lvs_Latn 55.67 65.02 63.69 67.42
norwegian
norwegian nynorsk (m:written)
nno_Latn nob_Latn 85.66 79.94 89.20 79.06
norwegian bokmål (m:written)
nob_Latn nob_Latn 83.81 82.90 83.82 84.14
sinitic
classical-middle-modern sinitic (o:traditional)
zho_Hant zho_Hans 89.82 86.80 89.02 86.39
cantonese
yue_Hant zho_Hans 89.45 86.46 88.71 87.64
classical-middle-modern sinitic (o:simplified)
zho_Hans zho_Hans 88.74 86.38 88.86 89.15
sotho-tswana (s.30)
northern sotho
nso_Latn nso_Latn 35.62 28.16 34.86 13.55
southern sotho
sot_Latn nso_Latn 32.55 32.31 39.93 19.08
southwestern shifted romance
portuguese (a:european)
por_Latn spa_Latn 88.13 89.10 88.10 87.74
galician
glg_Latn spa_Latn 86.99 89.00 86.93 87.83
spanish
spa_Latn spa_Latn 86.74 85.93 84.87 86.55
occitan
oci_Latn lij_Latn 84.12 74.80 78.53 62.56
@@
Table 20: Machine Translation

### v2

@@ Appendix E Evaluation results
anglic english eng_Latn eng_Latn 81.90 83.35 81.95 83.43
arabic standard arabic arb_Arab eng_Latn 65.57 73.83 65.57 73.85
najdi arabic ars_Arab eng_Latn 59.42 69.02 59.14 68.94
ta’izzi-adeni arabic acq_Arab eng_Latn 58.84 68.72 58.64 68.62
moroccan arabic ary_Arab eng_Latn 54.65 58.30 54.61 58.14
egyptian arabic arz_Arab eng_Latn 54.53 65.87 53.86 65.70
south levantine arabic ajp_Arab eng_Latn 54.09 64.13 53.42 63.81
north mesopotamian arabic acm_Arab eng_Latn 52.95 58.94 52.84 58.75
levantine arabic (a:north) apc_Arab eng_Latn 52.14 61.74 51.40 61.31
tunisian arabic aeb_Arab eng_Latn 47.54 50.18 47.42 50.20
common turkic north azerbaijani azj_Latn eng_Latn 59.76 73.17 59.20 73.17
central oghuz (m:spoken) tur_Latn eng_Latn 59.14 74.47 58.37 74.52
south azerbaijani azb_Arab eng_Latn 46.05 41.82 44.58 39.24
gallo-italian venetian vec_Latn eng_Latn 65.15 68.52 64.99 68.55
lombard lmo_Latn eng_Latn 59.38 56.39 59.34 56.16
ligurian lij_Latn eng_Latn 57.82 57.70 56.70 57.16
high german luxemburgish ltz_Latn eng_Latn 60.34 47.33 60.01 46.21
limburgan lim_Latn eng_Latn 50.80 59.88 50.31 59.75
italian romance italian ita_Latn eng_Latn 73.71 78.08 73.71 78.19
sicilian scn_Latn eng_Latn 62.69 56.17 62.66 55.82
kurdish central kurdish ckb_Arab eng_Latn 40.98 44.93 37.40 39.59
northern kurdish kmr_Latn eng_Latn 39.10 63.45 33.93 63.26
latvian latvian lvs_Latn eng_Latn 60.14 73.57 59.95 73.63
east latvian ltg_Latn eng_Latn 48.62 54.19 47.02 53.54
norwegian norwegian bokmål (m:written) nob_Latn eng_Latn 72.44 79.44 72.45 79.51
norwegian nynorsk (m:written) nno_Latn eng_Latn 68.08 71.16 68.10 71.06
sinitic classical-middle-modern sinitic (o:simplified) zho_Hans eng_Latn 68.52 72.61 68.54 72.57
classical-middle-modern sinitic (o:traditional) zho_Hant eng_Latn 61.72 64.89 61.48 64.49
cantonese yue_Hant eng_Latn 60.44 68.02 60.27 67.41
sotho-tswana (s.30) northern sotho nso_Latn eng_Latn 39.38 40.18 35.06 35.98
southern sotho sot_Latn eng_Latn 39.36 39.30 34.62 34.16
southwestern shifted romance spanish spa_Latn eng_Latn 75.13 79.00 75.15 79.09
portuguese (a:european) por_Latn eng_Latn 73.73 79.18 73.73 79.22
galician glg_Latn eng_Latn 73.39 78.48 73.39 78.55
occitan oci_Latn eng_Latn 68.48 63.09 68.47 62.96
@@
cluster variety target-code count Finetune (mBERT) Finetune (XLM-R) Zeroshot (mBERT) Zeroshot (XLM-R) ICL Mistral
anglic english (a:scotland) english–gbr 440 76.38 70.34 71.82 63.15 70.18
southern african english english–zaf 440 76.66 71.18 71.49 63.87 71.14
new zealand english english–nzl 440 76.71 71.39 71.22 63.69 70.95
australian english english–aus 440 75.66 70.89 71.20 62.28 69.23
southeast american english english–usa 440 77.26 71.50 71.17 63.71 71.76
irish english english–irl 440 75.52 70.73 70.92 62.15 70.64
philippine english english–phl 440 76.37 70.64 70.47 62.22 70.11
nigerian english english–nga 440 73.61 68.33 69.10 61.27 68.10
indian english (a:north) english–ind_n 440 74.62 68.03 68.84 61.25 68.99
kenyan english english–kenya 440 72.59 66.68 68.72 58.64 64.91
indian english (a:south) english–ind_s 440 71.93 66.88 66.49 60.36 65.13
arabic arabic (a:bahrain) arabic–bhr 921 77.52 72.11 53.25 53.28 55.33
arabic (a:jordan) arabic–jor 921 77.35 71.29 52.72 53.72 54.93
arabic (a:saudi-arabia) arabic–sau 921 77.88 72.11 52.72 53.24 55.66
algerian arabic arabic–dza 921 77.85 72.34 52.56 53.52 55.18
tunisian arabic arabic–tun 921 76.72 71.64 52.28 52.94 55.03
moroccan arabic arabic–mar 921 76.73 71.57 51.86 52.17 53.92
egyptian arabic arabic–egy 921 76.53 70.75 51.80 51.99 54.66
bengali bengali (a:west bengal) bengali–ind 113 68.62 73.27 32.30 36.39 50.06
bengali (a:dhaka) bengali–dhaka 113 67.37 74.24 31.79 35.52 51.73
korean seoul (m:spoken) korean–korn 276 10.15 31.91 7.26 19.62 65.79
korean (a:south-eastern, m:spoken) korean–kors 276 9.92 31.01 7.22 20.08 64.63
swahili swahili (a:tanzania) swahili–tanzania 472 63.54 62.30 38.24 39.38 55.70
swahili (a:kenya) swahili–kenya 472 72.25 70.53 37.97 41.59 58.71
@@
cluster variety dialect-code support precision (mBERT) precision (XLM-R) recall (mBERT) recall (XLM-R) F1 (mBERT) F1 (XLM-R)
anglic english (a:uk) english:en-gb 249 98.10 89.57 83.13 71.59 90.00 79.58
north american english english:en-us 349 93.27 88.14 83.38 82.09 88.05 85.01
arabic aleppo arabic:ale 200 59.50 58.50 59.50 48.35 59.50 52.94
algerian arabic arabic:alg 272 79.00 65.50 58.09 62.68 66.95 64.06
arabian peninsula arabic (a:yemen) arabic:san 177 60.50 55.50 68.36 56.63 64.19 56.06
egyptian arabic (a:alx) arabic:alx 192 70.50 74.50 73.44 66.82 71.94 70.45
egyptian arabic (a:asw) arabic:asw 221 56.00 52.00 50.68 45.02 53.21 48.26
egyptian arabic (a:cai) arabic:cai 130 35.50 36.50 54.62 72.28 43.03 48.50
egyptian arabic (a:kha) arabic:kha 244 63.50 55.50 52.05 44.05 57.21 49.12
fez. meknes arabic:fes 196 60.00 59.50 61.22 56.40 60.61 57.91
gilit mesopotamian arabic arabic:bag 203 57.50 47.50 56.65 49.48 57.07 48.47
gulf arabic (a:doh) arabic:doh 205 50.00 43.50 48.78 45.55 49.38 44.50
gulf arabic (a:jed) arabic:jed 196 58.00 46.00 59.18 40.89 58.59 43.29
gulf arabic (a:mus) arabic:mus 178 38.50 49.50 43.26 42.67 40.74 45.83
gulf arabic (a:riy) arabic:riy 311 62.00 56.50 39.87 37.92 48.53 45.38
levantine arabic (a:north-dam) arabic:dam 148 37.50 24.50 50.68 42.98 43.10 31.21
libyan arabic (a:ben) arabic:ben 238 56.50 47.50 47.48 52.78 51.60 50.00
north mesopotamian arabic (a:bas) arabic:bas 186 49.50 39.00 53.23 49.68 51.30 43.70
north mesopotamian arabic (a:mos) arabic:mos 188 71.50 70.00 76.06 69.31 73.71 69.65
rabat-casablanca arabic arabic:rab 153 50.00 40.00 65.36 60.61 56.66 48.19
sfax arabic:sfx 215 62.50 64.50 58.14 48.13 60.24 55.13
south levantine arabic (a:south-amm) arabic:amm 177 40.50 35.00 45.76 35.53 42.97 35.26
south levantine arabic (a:south-jer) arabic:jer 202 48.50 47.00 48.02 40.34 48.26 43.42
south levantine arabic (a:south-sal) arabic:sal 167 46.00 66.50 55.09 59.11 50.14 62.59
standard arabic arabic:msa 244 75.00 98.00 61.48 95.61 67.57 96.79
sunni beiruti arabic arabic:bei 192 58.00 52.50 60.42 68.18 59.18 59.32
tripolitanian arabic arabic:tri 201 66.00 60.00 65.67 60.30 65.84 60.15
tunisian arabic (a:tun) arabic:tun 164 52.50 37.00 64.02 56.49 57.69 44.71
greek cypriot greek (r:casual, m:written, i:other) greek:cg_other 81 74.14 84.48 53.09 56.32 61.87 67.59
cypriot greek (r:casual, m:written, i:twitter) greek:cg_twitter 36 51.11 44.44 63.89 68.97 56.79 54.05
modern greek (r:casual, m:written, i:twitter) greek:smg_twitter 94 89.83 100.00 56.38 53.15 69.28 69.41
high german central alemannic (a:be) swiss-dialects:be 389 72.89 51.58 71.21 62.42 72.04 56.48
central alemannic (a:bs) swiss-dialects:bs 340 73.50 57.83 75.88 61.14 74.67 59.44
central alemannic (a:lu) swiss-dialects:lu 335 72.91 65.13 75.52 59.47 74.19 62.17
central alemannic (a:zh) swiss-dialects:zh 359 78.84 73.33 75.77 63.73 77.27 68.19
sinitic mandarin chinese (a:mainland, o:simplified) mandarin_simplified:m 986 97.90 96.10 99.29 90.66 98.59 93.30
mandarin chinese (a:mainland, o:traditional, i:synthetic) mandarin_traditional:m 977 96.80 92.10 99.08 95.74 97.93 93.88
mandarin chinese (a:taiwan, o:simplified) mandarin_simplified:t 1014 99.30 90.10 97.93 95.85 98.61 92.89
mandarin chinese (a:taiwan, o:traditional, i:synthetic) mandarin_traditional:t 1023 99.10 95.90 96.87 92.39 97.97 94.11
southwestern shifted romance brazilian portuguese portuguese:pt-br 627 96.94 92.35 90.91 84.98 93.83 88.51
latin american spanish spanish:es-ar 207 81.06 9.25 88.89 91.30 84.79 16.80
portuguese (a:european) portuguese:pt-pt 349 91.45 83.64 70.49 63.92 79.61 72.46
portuguese (m:written) portuguese:pt 15 9.70 0.00 86.67 0.00 17.45 0.00
spanish spanish:es 290 74.21 74.53 81.38 47.69 77.63 58.16
spanish (a:europe) spanish:es-es 492 90.99 83.33 82.11 78.89 86.32 81.05
@@
arabic standard arabic arb_Arab arb_Arab 85.25 83.96 86.71 82.27
ta’izzi-adeni arabic acq_Arab arb_Arab 84.96 82.05 86.44 81.98
najdi arabic ars_Arab arb_Arab 84.80 84.39 87.41 83.33
north mesopotamian arabic acm_Arab arb_Arab 82.97 80.95 84.77 80.36
south levantine arabic ajp_Arab arb_Arab 81.82 80.16 84.16 79.05
levantine arabic (a:north) apc_Arab arb_Arab 81.59 80.15 83.76 79.88
egyptian arabic arz_Arab arb_Arab 81.02 76.38 84.43 81.03
tunisian arabic aeb_Arab arb_Arab 79.45 72.88 83.97 77.33
moroccan arabic ary_Arab arb_Arab 73.87 79.14 78.76 78.55
common turkic north azerbaijani azj_Latn azj_Latn 80.46 79.87 82.00 79.55
central oghuz (m:spoken) tur_Latn azj_Latn 79.10 84.41 80.61 79.51
south azerbaijani azb_Arab azj_Latn 65.90 67.08 69.71 68.37
gallo-italian venetian vec_Latn ita_Latn 76.72 70.68 75.07 74.28
lombard lmo_Latn ita_Latn 69.92 59.90 70.65 64.56
ligurian lij_Latn lij_Latn 66.81 63.42 74.03 57.78
high german luxemburgish ltz_Latn nld_Latn 74.74 58.50 77.86 64.83
limburgan lim_Latn nld_Latn 71.09 65.83 71.12 65.73
italian romance italian ita_Latn ita_Latn 87.67 84.92 86.68 85.83
sicilian scn_Latn ita_Latn 75.22 59.71 72.70 59.47
kurdish northern kurdish kmr_Latn ckb_Arab 33.23 68.21 10.45 5.71
central kurdish ckb_Arab ckb_Arab 13.10 19.37 16.86 12.38
latvian latvian lvs_Latn lvs_Latn 76.35 83.75 80.63 82.80
east latvian ltg_Latn lvs_Latn 55.67 65.02 63.69 67.42
norwegian norwegian nynorsk (m:written) nno_Latn nob_Latn 85.66 79.94 89.20 79.06
norwegian bokmål (m:written) nob_Latn nob_Latn 83.81 82.90 83.82 84.14
sinitic classical-middle-modern sinitic (o:traditional) zho_Hant zho_Hans 89.82 86.80 89.02 86.39
cantonese yue_Hant zho_Hans 89.45 86.46 88.71 87.64
classical-middle-modern sinitic (o:simplified) zho_Hans zho_Hans 88.74 86.38 88.86 89.15
sotho-tswana (s.30) northern sotho nso_Latn nso_Latn 35.62 28.16 34.86 13.55
southern sotho sot_Latn nso_Latn 32.55 32.31 39.93 19.08
southwestern shifted romance portuguese (a:european) por_Latn spa_Latn 88.13 89.10 88.10 87.74
galician glg_Latn spa_Latn 86.99 89.00 86.93 87.83
spanish spa_Latn spa_Latn 86.74 85.93 84.87 86.55
occitan oci_Latn lij_Latn 84.12 74.80 78.53 62.56
@@
Table 20: Zero-shot results for Machine Translation. We evaluate NLLB_600m and NLLB_1_3bn by translating each dialectal variety to English. For all languages without reference translations in English, we evaluate dialectal translations using the standard language’s translation as the reference. For varieties with parallel data like Yoruba, Turkish, Farsi, Tigrinya and Bengali we use the English reference provided in the dataset.